OpenStudy (anonymous):

Use the fact that the trigonometric functions are periodic to find the exact value of the expression. sin405 degrees tan405 degrees csc 350 sin 9pi ---- 4

OpenStudy (dumbcow):

sin(405) = sin(405-360) = sin(45) = sqrt2/2

OpenStudy (dumbcow):

tan(405) = tan(45) = 1

OpenStudy (anonymous):

I thought that the period for tan was 180

OpenStudy (dumbcow):

csc(350) = csc(-10) = -csc(10) = -1/sin(10) = -5.759

OpenStudy (dumbcow):

it is...so you go 2 periods but you still end up with reference angle of 45

OpenStudy (anonymous):

okay gotcha!

OpenStudy (anonymous):

please explain the csc

OpenStudy (dumbcow):

350 is not quite 1 period 350 -360 = -10 trig identity: csc(-theta) = -csc(theta)

OpenStudy (anonymous):

how do u get this: 1/sin(10) = -5.759

OpenStudy (dumbcow):

the last one is identical to the first except its in radians --> sin(9pi/4) = sqrt2/2

OpenStudy (dumbcow):

oh i used a calculator...if you have to have an exact expression you have to use double and half-angle formulas

OpenStudy (anonymous):

list all the trig identities please

OpenStudy (anonymous):

sin 9pi ---- 4 what about that problem?